# Evaluating LLMs on Sequential API Call Through Automated Test Generation

### v1

@@ V-A Limitation and Future Work
(1) Engineering effort for incorporating new scenarios: Our state machine currently requires manual modelling of each API’s documentation before it can generate programs for different scenarios. This process can be labor-intensive. Future work could reduce this effort by developing LLM-enabled middleware that automatically parses API documentation and translates it into structured state and transition models.
(2) More in-depth analysis: Due to space constraints, our analysis is necessarily limited in scope. Future work can extend our benchmark along several dimensions: (1) by evaluating a broader spectrum of large language models and prompting techniques, including advanced reasoning strategies such as chain-of-thought; and (2) by investigating methods to enhance the correctness and robustness of LLM-generated multi-function calls, particularly in complex or compositional scenarios.
@@ V-B Threat to Validity
External Validity. A potential threat to external validity arises from the possibility that StateGen’s effectiveness in generating diverse and challenging test cases may not generalize to different kinds of APIs. To mitigate this, we selected three substantially different domains—RESTful API calls (Session Service), tensor computations (Tensor Operation), and multimodal tool use (ElevenLabs MCP)—to introduce domain-level diversity and improve the generalizability of our findings.
@@
VI Conclusion
@@
In this paper, we introduced StateGen, an automated framework designed to evaluate the ability of large language models (LLMs) to generate code involving multiple API calls with complex interdependencies. Our approach begins with API sequence fuzzing using state machines, followed by a multi-agent process that translates code into natural language instructions. Leveraging this framework, we generated and manually curated 120 test cases across three diverse application scenarios, establishing a new benchmark, StateEval. Experimental results demonstrate that StateGen offers distinct advantages in generating sequential API calls with intricate dependencies. Evaluation of state-of-the-art LLMs on StateEval reveals that there remains substantial room for improvement in their tool usage capabilities. To support reproducibility and encourage further research, we will make both our codebase and the StateEval benchmark publicly available. We hope that our work will serve as a foundation for future research and development, ultimately leading to LLM systems with more robust and effective tool integration.

### v2

@@ V-A Limitation and Future Work
(1) Engineering effort for incorporating new scenarios: Our state machine currently requires manual modelling of each API’s documentation before it can generate programs for different scenarios. This process can be labor-intensive. Future work could reduce this effort by developing LLM-enabled middleware that automatically parses API documentation and translates it into structured state and transition models. While we acknowledge this limitation, our framework can still generate an extensive range of test cases within the existing three scenarios. This diversity stems from our fuzzer-like test generation engine. As the number of APIs in a program sequence increases, the number of diverse programs expands exponentially. Our engine can explore this large space by varying not only the API calls but also the program initial states and the API parameter combinations. This allows our current artifact to be used for extensive testing of subject LLMs.
(2) More in-depth analysis: Future work can extend our benchmark in three directions: (1) evaluate a broader range of models and prompting techniques, including advanced reasoning strategies (e.g., chain-of-thought); (2) develop automated root cause analysis to detect issues in generated code and enable self-reflection; and (3) improve the correctness and robustness of LLM‑generated sequences of function calls, especially in complex or compositional settings.
(3) Beyond correctness evaluation: While our work centers on code correctness, the framework’s core methodology—combining automata-based stateful generation with LLM-based transformation—is adaptable for evaluating other critical properties. For example, the fuzzing engine can be modified to generate malicious code or programs with vulnerabilities, effectively turning the framework into a tool for security testing. Likewise, it can assess privacy risks by having an LLM inject sensitive information into the generated code to test for potential data leaks.
@@ V-B Threat to Validity
External Validity. A potential threat to external validity arises from the possibility that StateGen’s effectiveness in generating diverse and challenging test cases may not generalize to different kinds of APIs. To mitigate this, we selected three substantially different domains—RESTful API calls (Session Service), tensor computations (Tensor Operation), and multimodal tool use (ElevenLabs MCP)—to introduce domain-level diversity and improve the generalizability of StateGen.
@@
Construct Validity. A potential threat to construct validity is whether the selected metrics truly capture the intended concepts of benchmark complexity and LLM proficiency. To address this, we evaluate from multiple perspectives. For benchmark complexity, we combine general metrics, such as instruction length and number of functions, with specialized metrics like Path Depth and Binding Count. For LLM proficiency, we complement the pass@1 with quantitative analyses of observed behaviors and case studies that examine underlying root causes.
VI Related Work
In this section, we review related work from three perspectives: (1) benchmarks for evaluating the correctness of generated programs by LLMs, (2) studies assessing LLMs’ ability to perform API calls, and (3) RESTful API fuzzing methods that involve sequential trace generation.
VI-A LLM Code Generation and Benchmark
LLMs have led to a significant performance leap in code synthesis, with models such as Gemini 2.5 [comanici2025gemini] and Qwen-Coder [hui2024qwen2] demonstrating strong performance on programming tasks. Despite the swift advancements, the evaluation of such coding capabilities remains fragmented across task granularity, complexity, and oracle quality. In particular, the evaluation of LLM code generation, at the early stage, focused on function-level code synthesis with executable unit test suites. HumanEval [chen2021codex, liu2023your] presented the widely-accepted pass@k metric using manually crafted Python problems, offering execution-based oracles for assessing function correctness. The MBPP [austin2021program] expanded its range to include approximately 1000 beginner Python tasks, each accompanied by three test cases, serving as a standard entry-level benchmark for evaluating the coding ability of LLMs.
In addition to function-level synthesis, practitioners constructed benchmarks that included libraries, APIs, and more stringent oracles to better reflect the complexity of real-world software development processes. DS-1000 [lai2023ds] collected 1000 programming problems associated with seven Python libraries (e.g., NumPy and Pandas) from StackOverflow and introduced reliable automatic evaluation by combining test suites and constraints on API usages or keywords. Meanwhile, HumanEval-XL [peng2024humaneval] expanded the benchmark and evaluation to a multilingual context, thereby investigating the cross-lingual generalization of code LLMs.
Building on function-level code synthesis, SWE-bench [jimenez2023swe] pioneered a repository-level assessment utilizing actual GitHub repositories and issues. That is, LLMs are required to create patches for projects that involve multiple files. Complementarily, BigCodeBench [zhuo2025bigcodebench] focuses on assessing the capabilities of code LLMs by utilizing functions from diverse libraries.
VI-B Evaluating LLMs with API Call
Substantial efforts have been made by practitioners to assess the capability of LLMs incorporating APIs under various scenarios [qin2023toolllm, openaiFC, mistralFC, llamaFC] with natural language instructions. Nevertheless, most existing approaches focus on building benchmarks or performing empirical studies by either manually collecting test cases or extracting them from online forums [li2023api, patil2024gorilla, basu2024nestful, huangmetatool, zhan2024injecagent, shen2024shortcutsbench, zhong2024can].
Early studies have focused on evaluating LLMs’ ability to select the appropriate tool and invoke it at the right time, leveraging metrics such as accuracy for evaluation [li2023api, qin2023toolllm, basu2024nestful]. For example, Qin et al. developed a benchmark for tool-augmented LLMs that assesses the model’s capability to choose the correct API and populate its parameters based on user queries [qin2023toolllm]. Although these studies often cover a wide range of domains collected in real-life, their tasks typically involve fulfilling specific requests rather than generating complete, executable software programs.
BFCL [patil2024gorilla] is another example that provides comprehensive evaluations of LLM performance on tasks involving API calls. However, the testing scenarios in BFCL still only involve a limited set of APIs, and the API calls typically lack interdependencies between their parameters.
VI-C Automated Testing of RESTful APIs
Automated generation of effective and diverse API call sequences remains a commonly studied yet challenging task. One representative research in this field is RESTful API testing [arcuri2019restful, atlidakis2019restler, viglianisi2020resttestgen, martin2021restest, wu2022combinatorial, kim2022automated, kim2023adaptive, deng2023nautilus, golmohammadi2023testing, wei2024demystifying], which generally follows a black-box approach and relies on standardized specifications such as OpenAPI [openapi]. This line of work typically adopts a black-box approach, as the source code is often unavailable in API testing. In such cases, Golmohammadi et al. identified the generation of meaningful API sequences as a primary challenge, which is complicated by inter-API dependencies and underspecified schemas [golmohammadi2023testing].
To address this challenge, techniques for API sequence generation largely fall into two categories. Top-down and bottom-up approaches. For the former, RESTTESTGEN [viglianisi2020resttestgen] is a representative case that first constructs an overall API call graph (e.g., an Operation Dependency Graph), and then instantiates API calls based on this graph. While providing a holistic view, the efficacy of such methods can be compromised by the quality of the initial graph, which may be derived from flawed or incomplete specifications.
Conversely, bottom-up strategies [atlidakis2019restler, liu2022morest] build sequences incrementally. For example, RESTler [atlidakis2019restler] extends sequences one call at a time using heuristics, while MOREST [liu2022morest] utilizes a dynamically updating RESTful-service Property Graph to capture detailed API behaviors and dependencies.
More recently, several studies have explored integrating LLMs into REST API test-case generation. In these approaches, LLMs are used to produce realistic test inputs, capture inter-parameter dependencies to improve test efficiency [kim2025llamaresttest, kim2025multi], or assist in identifying valid test oracles [alonso2025satori]. While these efforts show promising results, their objectives differ fundamentally from ours: they use LLMs to enhance software testing (i.e., AI for SE), whereas our work uses software engineering techniques to evaluate LLM capabilities (i.e., SE for AI).
VII Conclusion
@@
In this paper, we introduced StateGen, an automated framework designed to evaluate the ability of large language models (LLMs) to generate code involving multiple API calls with complex interdependencies. Our approach begins with API sequence fuzzing using state machines, followed by a multi-agent process that translates code into natural language instructions. Leveraging this framework, we generated and manually curated 120 test cases across three diverse application scenarios, establishing a new benchmark, StateEval. Experimental results demonstrate that StateGen offers distinct advantages in generating sequential API calls with intricate dependencies. Evaluation of state-of-the-art LLMs on StateEval reveals that there remains substantial room for improvement in their tool usage capabilities. We further identify three common root causes underlying LLM failures to generate programs that faithfully meet users’ requirements. To support reproducibility and encourage further research, we will make both our codebase and the StateEval benchmark publicly available. We hope that our work will serve as a foundation for future research and development, ultimately leading to LLM systems with more robust and effective tool integration.